OpenStudy (anonymous):

what r the intercepts of y=x^2 - 36

OpenStudy (anonymous):

for x intercerpts make your y value to 0 so 0=x^2-36 and then move the 36 to the left which makes it 36=x^2 then square root the value and you get x=6 that is the x intercept. make sure when u write the answer you write it as (6,0). and then for the y intercept make the x value a 0. so y=0^2 -36 which becomes y= -36. for y intercepts the answer should be written as (0,-36). there

OpenStudy (anonymous):

i got x intercept as x = -6 and also x=6

OpenStudy (anonymous):

and y equals -36

OpenStudy (anonymous):

also i have another question kind pf likle it if you can help?

OpenStudy (anonymous):

yea i forgot that x can be both -6 and 6 yes

OpenStudy (anonymous):

20x^2 + 13x - 15 = 0 find the real solutions of the equation

OpenStudy (anonymous):

ok for this u need to look at the b^2 -4ac= 0 equation

OpenStudy (anonymous):

if its more than 0 then you have 2 real solutions. if its less than 0 then u have no real solutions and its exactly 0 then u have one solution.

OpenStudy (anonymous):

so here the a= 20 b =13 c = -15

OpenStudy (anonymous):

did i confuse u but for this u need the quadratic formula. x =( -b+/- sqrt(b^2-4ac))/2a

OpenStudy (anonymous):

but when u plug in all numbers it has a negative under the square root.and we cant take a negative square root of a number which makes this equation with NO real solutions.